# Automated Federated Learning via Informed Pruning

### v1

@@ 3  Methodology
Exploration phase (Lines 2–3). The preliminary exploration phase aimed at understanding the relevance of each parameter (weight or bias) in view of loss improvement starts. For each client participating (in this study we select all the available clients), a local guidance matrix storing parameter deviations is computed.
@@
FL round with reduced client models (Lines 10–14). The standard algorithm FedAvg (McMahan et al., 2023) is used on the reduced framework. The pruned clients are trained. The server receives the local model updates and, upon aggregation, proceeds to update the global model with the FL aggregation strategy. Once updated, the global model is either ready for the next communication round or deemed ready for deployment if the convergence criteria are satisfied.
Accuracy-Improved Model Compression
A key aspect of AutoFLIP is its ability to maintain or potentially improve model accuracy during compression. Its iterative updating technique, powered by automated parameter importance evaluation, continuously refines the compression strategy, ensuring alignment with evolving data distributions and learning objectives. In non-IID data environments, AutoFLIP adapts its compression approach to suit each client’s dataset characteristics, preventing disproportionate impact on clients with less represented data features and safeguarding overall model accuracy across the network.
The global pruning mask, $PG_{\text{global}}$, is designed to systematically minimize the variance of the trained parameters across different clients, $\sigma^{2}_{\Delta W}=\text{Var}[\Delta W_{1},\Delta W_{2},\ldots,\Delta W_{K}]$. As shown in Figure 2, we prune the parameters that are characterized by scarce variability during the exploration phase and are therefore heavily influenced by their initialization value. Conversely, we continue to include in training the parameters that exhibit the highest deviations during the exploration phase. This indicates that their evolution was marked by high gradients, resulting in rapid exploration over similar loss landscapes (given that we are addressing similar prediction tasks).
By reducing variance through the preliminary exploration procedure, $PG_{\text{global}}$ crucially improves the convergence efficiency of the global model. This ensures that the global model evolves efficiently while aligning with the collective learning objectives of all participating clients.
@@
Communication efficiency
AutoFLIP enhances communication efficiency in FL by reducing model sizes transmitted between clients and the server, thus lowering bandwidth requirements for FL rounds. Its selective updating mechanism ensures only essential parameters, those significantly affecting model performance, are communicated. Integrating AutoFLIP into FL systems enhances faster training and inference times, lower energy consumption, and improved model scalability. The appendices  D,  E and  F provide a detailed analysis of how AutoFLIP accelerates inference and improves training efficiency, demonstrating its significant role in lowering computational costs and boosting FL’s overall efficiency.
@@ 4  Experiments
Inspired by Hahn et al. (2022), we benchmark AutoFLIP across established datasets to evaluate its robustness and adaptability in various Non-IID environments. We explore three distinct partitioning approaches for creating strongly Non-IID conditions: a Pathological Non-IID scenario, which involves 20 clients, each using data from two distinct classes, employing MNIST with a six-layer CNN (7,628,484 parameters) and CIFAR10 with EfficientNet-B3 (10,838,784 parameters), a Dirichlet-based Non-IID scenario, which utilizes the Dirichlet distribution to distribute data among 20 clients, with varying class counts per client, using CIFAR100 and a ResNet architecture (23,755,900 parameters), and a LEAF Non-IID scenario, which adopts the LEAF benchmark with FEMNIST and Shakespeare datasets. For FEMNIST, a CNN architecture with 13,180,734 parameters is used. For Shakespeare, we consider a two-layer LSTM model with 5,040,000 parameters. Further details on these scenarios are provided in Appendix B.
4.1  Experimental Setup
We evaluate AutoFLIP against FedAvg and different SOTA federated pruning strategies, including Random, L1, L2, Similarity, and BN mask, as described in Wu et al. (2023). We use 20 clients, a batch size of 350, and a learning rate of 0.0003 over 200 FL rounds involving 5 clients each. The setup includes a server momentum of 0.9, an SGD optimizer with weight decay, up to 150 exploration epochs, and a pruning threshold of 0.3. Data is split into 80-20 for training and testing, with model performance assessed by average prediction accuracy on test sets. We run 10 experiment repetitions to ensure statistical validity, measuring the compression rate to gauge model size reduction and its impact. We conducted experiments on two workstations, one with an Intel Core i5 and 16 GB RAM, and another with an Intel Xeon X5680, 128 GB of DDR4 RAM, and an NVIDIA TITAN X GPU.
4.2  Results
In Figure 3, we show the accuracy convergence profiles for the global model. The different federated pruning strategies are evaluated on both image recognition and text prediction tasks using five distinct datasets: MNIST, CIFAR10, CIFAR100, FEMNIST, and Shakespeare. Due to the varying complexities of each task, we use different model structures for different datasets. A plot depicting the evolution of the loss metric is presented in Appendix C.
@@ 4.2  Results
Here, AutoFLIP achieves an average client compression rate of x1.74. At each round, we remove on average 3244298 parameters for the six-layer CNN. For the EfficientNet-B3, we obtain an average compression rate of x2.1 with 5677458 deleted parameters. For the baselines we compare AutoFLIP to, we ensure that the number of parameters pruned matches the compression ratio of AutoFLIP, quantified as 42% for the six-layer CNN and 52.38% for EfficientNet-B3.
The first two subplots in Figure 3 show the evolution of global model accuracy during the FL rounds for the four-layer CNN with the MNIST dataset and for EfficientNet-B3 with the CIFAR10 dataset. In the case of the MNIST, the early rounds of FL show that AutoFLIP achieves slightly higher accuracy compared to both FedAvg and the other federated pruning strategies, among which RandomPruning emerges as the top performer. This indicates a faster convergence rate for our proposed method. However, the performance of the three baselines soon becomes comparable, with no clear superiority as the FL procedure progresses. We attribute this to the simplicity of the prediction tasks on the MNIST dataset compared to the excessive complexity of the four-layer CNN, which already possesses extremely good prediction capabilities that cannot be further enhanced by pruning. For the CIFAR10 dataset, we do not observe any advantage in using AutoFLIP over the other baselines. Surprisingly, all methods exhibit severe fluctuations in the accuracy convergence profiles up to FL round 100, after which they stabilize and become comparable.
Dirichlet-based Non-IID
For ResNet, AutoFLIP achieves an average compression rate for the clients of x1.58, with 8,720,520 parameters pruned on average out of 23,755,900 total parameters. Hence, we adjust the percentage of parameters to be pruned to 36.71% for the different baselines.
The third subplot in Figure 3 illustrates the evolution of the global model accuracy during the FL rounds for ResNet on CIFAR100. Here, AutoFLIP exhibits a performance enhancement throughout the considered training rounds. At round 200, it achieves an accuracy of 0.987, compared to 0.918 for FedAvg and 0.925 for RandomPruning. This enhancement signifies the robustness of AutoFLIP, showcasing its ability to maintain elevated performance levels, even when integrated with larger-complex neural networks and larger datasets.
LEAF Non-IID
In this scenario, AutoFLIP achieves an average compression rate of x1.8 for 5858104 client parameters pruned out of 13180734. Hence, we adjust the number of parameters to be pruned for the different baselines to 44%. As observed in the last two subplots of Figure 3 for the FEMNIST and Shakespeare datasets, AutoFLIP consistently outperforms the other pruning strategies by a significant margin. What stands out is the initial acceleration in convergence speed observed for AutoFLIP, firmly establishing it as a superior choice over FedAvg, RandomPruning and the other baselines in terms of maintaining limited computation and communication costs (see Appendices D and  E). Furthermore, this superiority persists throughout the entire FL training procedure. The final average accuracy values are 0.985 for AutoFLIP, 0.905 for FedAvg, and 0.935 for RandomPruning on the FEMNIST dataset. For the Shakespeare dataset, the values are 0.815, 0.783, and 0.738, respectively. Here, even L1 proves to be competitive, reaching a final accuracy equal to 0.802. However, it demonstrates inferior initial convergence. The remarkable performance of AutoFLIP in these tasks underscores its effectiveness in handling complex and realistic FL scenarios, thus emphasizing its applicability and robustness in real-world settings.
5  Conclusion and future perspectives
We presented AutoFLIP, an automated federated learning (FL) approach via informed pruning. Our method adjusts client DL model compression through informed pruning, leveraging information from a preliminary federated loss exploration phase. This enables efficient global model training and inference on clients with limited computational resources, such as network edge devices.
Extensive experiments across various non-IID scenarios show AutoFLIP’s ability to achieve high model accuracy while reducing computational and communication overhead. It improves convergence rates in federated settings with non-IID data distributions compared to FedAvg, with or without random pruning. AutoFLIP demonstrates adaptability and scalability across DL network types and multi-class datasets, showing notable improvements as task complexity increases. This underscores its potential for future research avenues, such as leveraging loss exploration for guiding more complex neural architecture search tasks.
While AutoFLIP has achieved promising results, we recognize that our analysis is not yet complete. Firstly, although to our knowledge we are the first to introduce automated informed pruning in an FL framework, we plan to conduct extensive comparisons with other SOTA pruning methods for FL. Additionally, validating AutoFLIP across various real-world domains such as healthcare, Internet of Things, and mobile computing could further demonstrate its effectiveness across different federated environments and data modalities.
6  Broader Impact Statement
After careful reflection, the authors have determined that this work presents no notable negative impacts on society or the environment. On the contrary, by introducing an automatic informed pruning strategy for federated learning, the method presented in this study offers a more sustainable and efficient utilization of edge devices in federated learning tasks.
Acknowledgements
Appendix A Ablation Study on $T_{p}$
Figure 4: Ablation on $T_{p}$ for MINST/Non-IID based on average accuracy (top) and loss (bottom).
Appendix B Partitioning approaches
Pathological Non-IID
@@ Appendix B Partitioning approaches
This experimental configuration is delineated by each client possessing data exclusively from two distinct classes within a broader multi-class dataset. Figure 6 illustrates this ”pathological” data partitioning scenario within the CIFAR10 dataset across 20 clients. For our experiments, we select the MNIST dataset (Deng, 2012) with a six-layer CNN (7628484 parameters) and the CIFAR10 dataset (Krizhevsky, 2012) with EfficientNet-B3 architecture (10838784 parameters), following the guidelines in (McMahan et al., 2023) and (Tingting et al., 2023).
@@
This advanced experimental setup, as introduced by Hsu et al. (2019), utilizes the Dirichlet distribution, modulated by a concentration parameter $\alpha$. Let $\boldsymbol{p}=(p_{1},p_{2},...,p_{N})$ be the class distribution for a given client, where $N$ is the number of classes. The Dirichlet distribution is defined as $\boldsymbol{p}\sim\text{Dir}(\alpha\cdot\boldsymbol{1}_{N})$, where “Dir” denotes the Dirichlet distribution, $\alpha$ is the concentration parameter, and $\boldsymbol{1}_{N}$ is a N-dimensional vector of ones. In this context, a low value of $\alpha$, or $\alpha\to 0$, leads to distributions where most of the probability mass is concentrated on a single class, thereby indicating that each client’s data is restricted to a single class. Conversely, as $\alpha\to\infty$, $\boldsymbol{p}$ approaches a uniform distribution, ensuring that the samples are evenly split across all clients. Figure 7 illustrates this “Dirichlet-based Non-IID” data partitioning scenario within the CIFAR100 dataset across 20 clients, with individual colors denoting separate classes.
@@
To address the complexities of larger datasets, we have extended our evaluation to include CIFAR100 (Krizhevsky, 2012) with a $\alpha=100$, employing ResNet (23755900 parameters) (He et al., 2016) in alignment with the methodology proposed in (Hahn et al., 2022).
@@
Utilizing the popular LEAF benchmark for FL (Caldas et al., 2019), we selected the FEMNIST and Shakespeare datasets to simulate closer real-world FL scenarios, with each dataset designed for specific tasks. The FEMNIST dataset is defined for a multi-class classification challenge involving 62 distinct classes. Conversely, the Shakespeare dataset is tailored for a next-character prediction task, requiring models to predict the subsequent character from a sequence of 80 characters, thereby testing the model capabilities in sequential data processing and language modeling. The incorporation of the next-character prediction task allows for a comprehensive assessment of AutoFLIP adaptability and performance across diverse task types and deep neural network architectures, such as Long Short-Term Memory (LSTM) networks.
In our experimental setup, we employed the FEMNIST-CNN architecture, as delineated in (Caldas et al., 2019), for the FEMNIST dataset. For the Shakespeare dataset, we utilized a two-layer (LSTM) (5040000 parameters) model, in accordance with the specifications provided in (McMahan et al., 2023).
Appendix C Loss plots

### v2

@@ 4.1 Federated Loss Exploration
Exploration phase (Lines 2–3).
The preliminary exploration phase aimed at understanding the relevance of each parameter (weight or bias) in view of loss improvement starts. For each client participating (in this study we select all the available clients), a local guidance matrix storing parameter deviations is computed.
@@
FL round with reduced client models (Lines 10–14). The standard algorithm FedAvg [38] is usedin this framework but AutoFLIP can be applied to other SotA FL aggregation algorithm. The pruned clients are trained. The server receives the local model updates and, upon aggregation, proceeds to update the global model with the FL aggregation strategy. Once updated, the global model is either ready for the next communication round or deemed ready for deployment if the convergence criteria are satisfied.
4.2 Robustness and Efficiency of AutoFLIP
Referring to [12, 11, 57, 60], we base our convergence guarantees on a federated stochastic aggregation scheme. The authors’ assumptions on Lipschitz smoothness, convexity of local loss functions, unbiased gradient estimators, finite client answering times, and specific client aggregation weights form the theoretical backbone of AutoFLIP. These conditions ensure that the learning process remains stable and converges efficiently even in the presence of non-IID data distributions. With AutoFLIP, at each round, each client experiences the same pruning strategy with the pruning mask $PG_{\text{global}}$, resulting in a substantial decrease in the variance ( $\sigma^{2}_{\Delta W}$) previously defined in Section 3.2 of weight updates for the global model. This uniform pruning strategy minimizes discrepancies in weight adjustments across clients by focusing updates on critical weights identified during the federated loss exploration phase. The reduction in variance helps to alleviate the bias caused by the non-IID setting, as shown in the work of [65], thus promoting better global convergence.
@@ 4.2 Robustness and Efficiency of AutoFLIP
Furthermore, [59, 45, 53, 24] provide a theoretical foundation for which pruned NNs can effectively learn signals. They demonstrate that pruning preserves the signal’s magnitude in features and reduces noise, leading to improved generalization. These studies highlight that pruning, when done correctly, does not degrade the model’s capacity to learn but rather focuses the learning on more relevant features. By focusing on parameters with significant contributions to the loss topology, AutoFLIP ensures that the essential features are retained. As illustrated in Figure 2 for different NN, the parameters in $G_{global}$ with minimal variability during the federated loss exploration phase are pruned, while those exhibiting high deviations are retained. Note that higher frequencies are recorded for smaller deviation values, indicating that many parameters, according to our pruning strategy, are non important. The high density of those parameters lead to the weights of entire channels or layers being set to zero. In such cases, we delete the entire substructure of the NN, resulting in a structured network reduction. This leads to an improvement in computing efficiency as detteiled in appendix  E.
AutoFLIP enhances communication efficiency in FL by reducing model sizes transmitted between clients and the server, thus lowering bandwidth requirements for FL rounds. Its selective updating mechanism ensures only essential parameters, those significantly affecting global model convergence, are communicated. The appendices  F,  G and  H provide a detailed analysis of how AutoFLIP accelerates inference and improves training efficiency, demonstrating its significant role in lowering computational costs and boosting FL’s overall applicability.
@@ 5 Experiments
Inspired by [16], we benchmark AutoFLIP across established datasets to evaluate its robustness in various non-IID environments. We explore three distinct partitioning approaches for creating strongly non-IID conditions: a Pathological non-IID scenario, which involves clients using data from two distinct classes, employing MNIST with a six-layer CNN (7,628,484 parameters) and CIFAR10 with EfficientNet-B3 (10,838,784 parameters), a Dirichlet-based non-IID scenario, which utilizes the Dirichlet distribution to distribute data among clients, with varying class counts per client, using CIFAR100 with ResNet (23,755,900 parameters), and a LEAF non-IID scenario, which adopts the LEAF benchmark [4] with FEMNIST and Shakespeare datasets. For FEMNIST, a CNN architecture with 13,180,734 parameters is used. For Shakespeare, we consider a two-layer LSTM model with 5,040,000 parameters. Further details on these scenarios are provided in Appendix C.
5.1 Experimental Setup and Results
We evaluate AutoFLIP against both FedAvg without any model compression and with SotA algorithm e.g PruneFL [26], and FL-pruning with various parameter selection criteria: Random, L1, L2, Similarity, and BN mask, as described in [58]. The experimental setup involves $C=20$ (for LEAF non-IID scenario we employ $C=730$ for Shakespeare and $C=660$ for FEMNIST) clients, a batch size $B=350$, a local update epochs $E=5$, and a learning rate $\eta=0.0003$ over $200$ total rounds $R$ with $K=5$ (for LEAF non-IID scenario $K=20$) clients selected per round. We incorporate a server momentum of $0.9$ and use an SGD optimizer with weight decay. The exploration phase consists of up to $E_{\text{exp}}=150$ epochs, and the pruning threshold is set to $T_{\text{p}}=0.3$. Data is divided into 80% for training and 20% for testing, with global model performance assessed by the average prediction accuracy on the test sets. To ensure statistical validity, each experiment is repeated 10 times. We measure the compression rate to evaluate model size reduction and its impact. Experiments were conducted with an Intel Xeon X5680, 128 GB of DDR4 RAM, and an NVIDIA TITAN X GPU.
@@ 5.1 Experimental Setup and Results
Pathological non-IID) Here, AutoFLIP achieves an average client compression rate of x1.74. At each round, we remove on average 3244298 parameters of the six-layer CNN for each participant client. For the EfficientNet-B3, we obtain an average compression rate of x2.1 with 5677458 deleted parameters. For a fair comparison with the baselines, we ensure that the number of parameters pruned matches the compression ratio of AutoFLIP, quantified as 42% for the six-layer CNN and 52.38% for EfficientNet-B3.
The first two subplots in Figure 3 show the evolution of global model accuracy during the FL rounds for the four-layer CNN with the MNIST dataset and for EfficientNet-B3 with the CIFAR10 dataset. Refer to Appendix D for the evolution of the loss metric. In the case of the MNIST, the early rounds of FL show that AutoFLIP achieves slightly higher accuracy compared to both FedAvg and the other FL pruning strategies, among which RandomPruning emerges as the top performer. This indicates a faster convergence rate for our proposed method. However, the performance of the baselines soon becomes comparable, with no clear superiority as the FL procedure progresses. We attribute this to the simplicity of the prediction tasks on the MNIST dataset compared to the excessive complexity of the four-layer CNN, which already possesses extremely good prediction capabilities that cannot be further enhanced by pruning. For the CIFAR10 dataset, we do not observe any advantage in using AutoFLIP over the other baselines. Surprisingly, all methods exhibit severe fluctuations in the accuracy convergence profiles up to FL round 100, after which they stabilize and become comparable.
Dirichlet-based non-IID) For ResNet, AutoFLIP achieves an average compression rate for the clients of x1.58, with 8,720,520 parameters pruned on average out of 23,755,900 total parameters. Hence, we adjust the percentage of parameters to be pruned to 36.71% for the different baselines.
The third subplot in Figure 3 illustrates the evolution of the global model accuracy during the FL rounds for ResNet on CIFAR100. Here, AutoFLIP exhibits a performance enhancement throughout the considered training rounds. At round 200, it achieves an accuracy of 0.987, compared to 0.918 for FedAvg and 0.925 for PruneFL. This enhancement signifies the robustness of AutoFLIP, showcasing its ability to maintain elevated performance levels when integrated with larger-complex neural networks and larger datasets.
LEAF non-IID) In this scenario, AutoFLIP achieves an average compression rate of x1.8 for 5858104 client parameters pruned out of 13180734. Hence, we adjust the number of parameters to be pruned for the different baselines to 44%. As observed in the last two subplots of Figure 3 for the FEMNIST and Shakespeare datasets, AutoFLIP consistently outperforms the other pruning strategies by a significant margin.
What stands out is the initial acceleration in convergence speed observed for AutoFLIP, firmly establishing it as a superior choice overFedAvg and the other FL baselines. Furthermore, this superiority persists throughout the entire FL training procedure. The final average accuracy values are 0.985 for AutoFLIP, 0.905 for FedAvg, and 0.935 for RandomPruning on the FEMNIST dataset. For the Shakespeare dataset, the values are 0.815, 0.783, and 0.738, respectively. Here, even L1 proves to be competitive, reaching a final accuracy equal to 0.802. However, it demonstrates inferior initial convergence.
6 Conclusion and Limitations
We introduced AutoFLIP, an innovative federated learning (FL) approach that employs informed pruning to optimize deep learning (DL) models on clients with limited computational resources. Through extensive experiments in various non-IID scenarios, AutoFLIP has demonstrated its ability to achieve better accuracy and significantly reduce computational and communication overheads. It enhances global convergence in federated settings and shows remarkable adaptability and scalability across diverse DL model architectures and multi-class datasets, particularly as the complexity of tasks increases.
Limitations. AutoFLIP shows promise but has limitations. It is primarily tested in the popular efficient single-server setting, not accounting for multi-server or hierarchical environments with diverse client capabilities and model structure. Our tests also assume standard conditions without data label noise.
Future Research Directions. AutoFLIP underscores its potential for future research avenues, such as leveraging loss exploration for guiding complex Neural Architecture Search (NAS) tasks. Enhancements will focus on refining AutoFLIP’s dynamic and adaptive pruning to better client personalization. We aim to perform comparison analysis with other strategies from other domain such us like NAS or Client Dropout. Further, the impact on data privacy and defense against adversarial clients during the federated loss exploration phase has to assessed. Research will also explore the extension of AutoFLIP to more complex DL architectures and its integration into real-world applications across various domains such as healthcare and mobile computing.
Broader Impact. AutoFLIP enhances sustainability and efficiency in FL, reducing the energy footprint of training deep learning models. Its utility in sensitive sectors like healthcare and finance emphasizes its societal importance. However, deploying AutoFLIP requires careful consideration of ethical issues, including data privacy and biases. Proactive management and regulation are crucial to ensure its positive societal impact and responsible integration into critical fields.
Appendix A Ablation Study on $C_{\text{exp}}$ and $E_{\text{exp}}$
We perform an ablation study to assess the sensitivitiy of our method to the parameters $C_{\text{exp}}$ and $E_{\text{exp}}$. The number of explorer clients $C_{exp}$ influences the comprehensiveness of $G_{\text{global}}$ in capturing the intricacies of the clients loss landscapes. The depth of the exploration phase, quantified by the number of exploration epochs $E_{exp}$ influences loss function surface understanding and $G_{\text{global}}$’s knowledge depth.
In particular, we check how the average accuracy and loss for the global model predictions vary for $C_{\text{exp}}\in\{0.25,0.5,0.75,2.0\}$ and for $E_{\text{exp}}\in\{150,300,500,750,1000\}$. We do this for datasets FEMINST the LEAF non-IID scenario 4. It is possible to observe in Figure 4, the x-axis represents the $E_{exp}$ parameter. Each distinct plot corresponds to a different $C_{exp}$, with the highest accuracy achieved distinctly highlighted. In particular, our findings reveal that even a conservative value of $C_{exp}$ can boost the accuracy. The influence of $C_{exp}$ is substantial, with higher counts of explorer clients resulting in improved initial accuracy, indicative of a more robust $G_{\text{global}}$ at the outset of the learning process. A discernible trend suggests that increasing the $E_{exp}$ value generally leads to an improvement in accuracy. However, at higher values, the increase is reduced, indicating a saturation point beyond which additional exploration epochs no longer improve accuracy.
Figure 4: Ablation on the total number of clients for exploration $C_{\text{exp}}$ and the total number of exploration epochs $E_{\text{exp}}$ for FEMNIST/pathological non-IID data, based on average accuracy.
Appendix B Ablation Study on $T_{p}$
Figure 5: Ablation on $T_{p}$ for MNIST/non-IID based on average accuracy (top) and loss (bottom).
Appendix C Partitioning approaches
@@
This experimental configuration is delineated by each client possessing data exclusively from two distinct classes within a broader multi-class dataset. Figure 7 illustrates this "pathological" data partitioning scenario within the CIFAR10 dataset across 20 clients. For our experiments, we select the MNIST dataset [6] with a six-layer CNN (7628484 parameters) and the CIFAR10 dataset [28] with EfficientNet-B3 architecture (10838784 parameters), following the guidelines in [38] and [52].
@@
This advanced experimental setup, as introduced by [22], utilizes the Dirichlet distribution, modulated by a concentration parameter $\alpha$. Let $\boldsymbol{p}=(p_{1},p_{2},...,p_{N})$ be the class distribution for a given client, where $N$ is the number of classes. The Dirichlet distribution is defined as $\boldsymbol{p}\sim\text{Dir}(\alpha\cdot\boldsymbol{1}_{N})$, where “Dir” denotes the Dirichlet distribution, $\alpha$ is the concentration parameter, and $\boldsymbol{1}_{N}$ is a N-dimensional vector of ones. In this context, a low value of $\alpha$, or $\alpha\to 0$, leads to distributions where most of the probability mass is concentrated on a single class, thereby indicating that each client’s data is restricted to a single class. Conversely, as $\alpha\to\infty$, $\boldsymbol{p}$ approaches a uniform distribution, ensuring that the samples are evenly split across all clients. Figure 8 illustrates this “Dirichlet-based non-IID” data partitioning scenario within the CIFAR100 dataset across 20 clients, with individual colors denoting separate classes.
@@
To address the complexities of larger datasets, we have extended our evaluation to include CIFAR100 [28] with a $\alpha=100$, employing ResNet (23755900 parameters) [19] in alignment with the methodology proposed in [16].
@@
Utilizing the popular LEAF benchmark for FL [4], we selected the FEMNIST and Shakespeare datasets to simulate closer real-world FL scenarios, with each dataset designed for specific tasks. The FEMNIST dataset is defined for a multi-class classification challenge involving 62 distinct classes. Conversely, the Shakespeare dataset is tailored for a next-character prediction task, requiring models to predict the subsequent character from a sequence of 80 characters, thereby testing the model capabilities in sequential data processing and language modeling. The incorporation of the next-character prediction task allows for a comprehensive assessment of AutoFLIP adaptability and performance across diverse task types and deep neural network architectures, such as Long Short-Term Memory (LSTM) networks.
In our experimental setup, we employed the FEMNIST-CNN architecture, as delineated in [4], for the FEMNIST dataset. For the Shakespeare dataset, we utilized a two-layer (LSTM) (5040000 parameters) model, in accordance with the specifications provided in [38].
Appendix D Loss plots